Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 5,200 units of product were as follows:

Standard Costs Actual Costs
Direct materials 6,800 lb. at $4.60 6,700 lb. at $4.40
Direct labor 1,300 hrs. at $18.10 1,330 hrs. at $18.50
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 1,360 direct
labor hrs.:
Variable cost, $3.20 $4,120 variable cost
Fixed cost, $5.10 $6,936 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials price variance $    Favorable/Unfavorable
Direct materials quantity variance Favorable/Unfavorable
Total direct materials cost variance $ Favorable/Unfavorable

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct labor rate variance    $    Favorable/Unfavorable
Direct labor time variance Favorable/Unfavorable
Total direct labor cost variance $ Favorable/Unfavorable

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $    Favorable/Unfavorable
Fixed factory overhead volume variance Favorable/Unfavorable
Total factory overhead cost variance $ Favorable/Unfavorable
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
6700 $4.40 $29,480.00 6700 $4.60 $30,820.00 6800 $4.60 $31,280.00
$1,340.00 F $460.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance -$1,340.00 F
Direct material quantity variance -$460.00 F
Direct material cost variance -$1,800.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
1330 $18.50 $24,605.00 1330 $18.10 $24,073.00 1300 $18.10 $23,530.00
$532.00 U $543.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $532.00 U
Direct Labor Efficiency variance $543.00 U
Direct labor cost variance $1,075.00 U

Solution c:

Variable factory overhead controllable variance = Standard cost of variable overhead - Actual cost of variable overhead

= (5200*0.25*$3.20) - $4,120 = -$40 F

Fixed factory overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= (5200*0.25*$5.10) - $6,936 = $306 U

Total factory overhead cost variance = $40 F + $306 U = $266 U